if a cyclist moving with a speed of 4.9m/s on a leval road can take a sharp circular turn of radius  4m , then coefficient of friction between the cycle tyres and road is?

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

If the cyclist is able to take the turn, the fricitonal force must be large enough to provide the centripetal acceleration.

Hencem\dfrac{v^2}{R}=\mu mg
\implies \mu=\dfrac{v^2}{gR}=0.61